OpenStudy (anonymous):

Express the complex number in trigonometric form. -3 + 3sqroot(3i)

zepdrix (zepdrix):

This is the expression? With the i under the root? \[\Large\bf\sf -3+3\sqrt{3i}\]

OpenStudy (anonymous):

correct

zepdrix (zepdrix):

Hmm I dunno :(

OpenStudy (gorv):

it should be\[-3+3\sqrt{3} *i\]

OpenStudy (gorv):

similar to standard form x+y i here x=r*cos(theta) y=r*sin(theta) and \[r=\sqrt{x^{2}+y ^{2}}\]

zepdrix (zepdrix):

Yes, \[\Large\bf\sf -3+\left(3\sqrt{3}\right)\mathcal i\]makes much more sense as it has a special angle representation in polar. But I verified it with the asker, he seems to think the i is under the root. \(\Large\bf\sf -3+3\sqrt{3i} \) `does` have a polar representiation, I'm just not sure how to find it.